OpenStudy (anonymous):

3(d-8)-5=9(d+2)+1 Can anyone help me? This is an Algebra 1 equation, Solving variables on each side! Thank you!

jhonyy9 (jhonyy9):

first elimine these parantheses 3d-24-5=9d+18+1 3d-29=9d+19 add to both sides 29 3d-29+29=9d+19+29 3d=9d+48 add to both sides -9d 3d-9d=9d-9d+48 -6d=48 divide both sides by -6 d=48/(-6) d=-8